OpenStudy (anonymous):

A manufacturer wants to design an open box with a square base and a surface area of 108in(squared). What dimensions will max volume? max volume?

OpenStudy (anonymous):

each side is 54 i think becuase 108/2...i got that from A=BH

OpenStudy (anonymous):

its surface area.. not area of the box.

OpenStudy (anonymous):

wait is that 108 from one side?

OpenStudy (anonymous):

oh its the...i got it..lol

OpenStudy (anonymous):

:D

OpenStudy (anonymous):

Yay jimmy :) REP!

OpenStudy (anonymous):

Yes please HELP ! :) diarmuid ! and JIMMMY !

OpenStudy (anonymous):

the maximum area of a rectangle is a square so as there are 5 square sides each will have area 108/5 = 21.6 and side sqrt 21.6 so maxm volume will be (sqrt21.6)^3

OpenStudy (anonymous):

no. only base is a square .. the sides are not

OpenStudy (anonymous):

sides have to be square also maximum volume = 100.4 cu ins

OpenStudy (anonymous):

you cant have rectangle sides and a square base

OpenStudy (anonymous):

height of the sides are not the same as "x" when x is the length and width of the base

OpenStudy (anonymous):

Got it thanks guys !

OpenStudy (anonymous):

x is the base y is the height so 108 = x^2 + 4 xy implies y = (-x^2 +108)/4x also maximise yx^2 (volume) substitue y so formula becomes (-x^3)/(4 +27x) so max (-x^3)/4 +27x draw for sanity ant it has a local max at x = +6 or so differentiate -3/4x^2 +27 = 0 so x = +-6 x = +6 108 = x^2 +4xy 72 = 4(6)y 72/24 = y y = 3 so box dimensions are 6*6*3

OpenStudy (anonymous):

AWESOME !

OpenStudy (anonymous):

No worries. Is this a college question?

OpenStudy (anonymous):

yup - good answer - i must have had a mental block on this one!